OpenStudy (anonymous):

whats the integral of (e^x-2) dx?

OpenStudy (anonymous):

So what is the derivative of e^x-2? Just go backwards. What would you take the derivative to get e^x-2? Also, quick question. Do you mean e^(x-2) or e^(x) -2? Both are completely different.

OpenStudy (anonymous):

i meant e^(x) -2

OpenStudy (anonymous):

e^2x?

OpenStudy (anonymous):

what no

OpenStudy (anonymous):

nevermind is it, e^x -2x?

OpenStudy (anonymous):

So are you saying that the equation is \[e^x -2x\] then?

OpenStudy (anonymous):

yeah

OpenStudy (sgadi):

\[\int {(e^x-2)dx}=e^x-2x+c\] where c is constant.

OpenStudy (anonymous):

that seems too simply...

OpenStudy (anonymous):

thanks for the help

OpenStudy (anonymous):

Oh sorry, I misunderstood what you meant when you said "nevermind is it, e^x -2x?" I thought that was what you were taking the integral of.

OpenStudy (anonymous):

its fine, i worded it wrong so no worries

OpenStudy (anonymous):

i've just spent an hour integrating using substitution so when i got an easy problem i got really confused...